what volume of oxygen will be required to oxidise 1.4 cubic dm of carbon(iv)oxide and what is the volume of the product ,all volumes measured at s.t.p


1
Expert's answer
2021-10-26T05:38:41-0400

Solution:

carbon(II) oxide: CO

oxygen gas: O2

carbon(IV) oxide: CO2


Balanced chemical equation:

2CO(g) + O2(g) → 2CO2(g)

According to stoichiometry:

Moles of CO = 2 × Moles of O2 = Moles of CO2

n(CO) = 2 × n(O2) = n(CO2)


n = V / Vm

At STP, one mole of any gas occupies a volume of 22.4 dm3 (Vm).

Hence,

V(CO) = 2 × V(O2) = V(CO2)

V(CO2) = V(CO) = 1.4 dm3

V(O2) = V(CO) / 2 = (1.4 dm3) / 2 = 0.7 dm3


Answer:

V(CO2) = V(CO) = 1.4 dm3

V(O2) = 0.7 dm3